# TPP-SD: Accelerating Transformer Point Process Sampling with Speculative Decoding

### v1

@@ D.2 Architectural Differences across Encoders
We incorporate the history encoders from THP, SAHP, and AttNHP into our CDF-based Transformer TPPs designed for TPP-SD. The architectural and implementation differences among these encoders may explain the varying sampling speed performance observed in Table 1 and Table 2.
Following the notation in Section 4.2, the temporal encoding $\mathbf{z}(t_{i})\in\mathbb{R}^{D}$ for each encoder differs slightly:
@@
The attention mechanisms differ across encoder architectures. For clarity, we present the single-head attention formulation for each model. The layer-$l$ history embedding of event $(t_{i},k_{i})$ is computed as:
@@
where $f(\cdot,\cdot)$ is a Gaussian kernel $f(\mathbf{q}^{(l)}(t_{j}),\mathbf{k}^{(l)}(t_{i}))=\exp(\frac{1}{\sqrt{D}}\mathbf{q}^{(l)}(t_{i})^{\top}\mathbf{k}^{(l)}(t_{j}))\in\mathbb{R}$. The key distinction lies in how query, key, and value vectors are derived at layer-$l$. In THP and SAHP, $\mathbf{q}^{(l)}(t_{i})=\mathbf{Q}^{(l)}\mathbf{h}^{(l-1)},\mathbf{k}^{(l)}(t_{i})=\mathbf{K}^{(l)}\mathbf{h}^{(l-1)},\mathbf{v}^{(l)}(t_{i})=\mathbf{V}^{(l)}\mathbf{h}^{(l-1)}$ are obtained through linear projection, whereas in AttNHP:

### v2

@@ D.2 Architectural Differences across Encoders
We incorporate the history encoders from THP, SAHP, and AttNHP into our CDF-based Transformer TPPs designed for TPP-SD. The architectural and implementation differences among these encoders may explain the varying sampling speed performance observed in Table˜1 and Table˜2.
Following the notation in Section˜4.2, the temporal encoding $\mathbf{z}(t_{i})\in\mathbb{R}^{D}$ for each encoder differs slightly:
@@
The attention mechanisms differ across encoder architectures. For clarity, we present the single-head attention formulation for each model. The $l$ history embedding of event $(t_{i},k_{i})$ is computed as:
@@
where $f(\cdot,\cdot)$ is a Gaussian kernel $f(\mathbf{q}^{(l)}(t_{j}),\mathbf{k}^{(l)}(t_{i}))=\exp(\frac{1}{\sqrt{D}}\mathbf{q}^{(l)}(t_{i})^{\top}\mathbf{k}^{(l)}(t_{j}))\in\mathbb{R}$. The key distinction lies in how query, key, and value vectors are derived at $l$. In THP and SAHP, $\mathbf{q}^{(l)}(t_{i})=\mathbf{Q}^{(l)}\mathbf{h}^{(l-1)},\mathbf{k}^{(l)}(t_{i})=\mathbf{K}^{(l)}\mathbf{h}^{(l-1)},\mathbf{v}^{(l)}(t_{i})=\mathbf{V}^{(l)}\mathbf{h}^{(l-1)}$ are obtained through linear projection, whereas in AttNHP: